I havae performed 1241 glucoses for 1 month. 289 glucoses were audited for errors. 12 errors were found out of the 289 glucoses audited. What is the percentage of error for thel month? would it be 1241x289/12
Answer by checkley77(12569)

12/289=x/1241
289x=12*1241
289x=14,892
x=14,892/289
x=51.529 errors in the total of 1241 tests.
51.529/1241=.04 or 4% error rate.